# Estimating LLM Uncertainty with Logits

### v2

@@ Abstract
In recent years, Large Language Models (LLMs) have seen remarkable advancements and have been extensively integrated across various fields. Despite their progress, LLMs are prone to hallucinations, producing responses that may not be reliable if the models lack sufficient grounding knowledge. To mitigate this issue, methods for estimating uncertainty have been adopted, with a focus on critical tokens as indicators of reliability. However, probability-based approaches have shown limitations in assessing token-level reliability due to the loss of evidence strength information acquired during training. In this paper, we introduce Logits-induced Token Uncertainty (LogU), a novel framework designed to estimate token-specific uncertainty in LLMs in real time, without the need for multiple sampling rounds. By leveraging evidence modeling for the implementation of LogU, we utilize the derived uncertainty measures to steer downstream tasks. Our experimental findings highlight the substantial effectiveness and potential of LogU, marking a significant advance in addressing the challenge of model hallucinations.
Machine Learning, ICML
@@ 1 Introduction
Over the past few years, Large Language Models (LLMs) have developed rapidly and LLM-driven systems are deployed across various domains. Despite their remarkable performance, LLMs remain prone to hallucinations (Banerjee et al., 2024), which causes them to generate unreliable responses when the models lack the corresponding knowledge. Hallucinations critically undermine the reliability of LLMs, particularly in professional applications such as medical and legal consultations (Shah, 2024; Dahl et al., 2024). Hallucinations have been considered as a major barrier to the broader deployment of LLMs (Huang et al., 2023; Liu et al., 2024; Perković et al., 2024; Zhou et al., 2024).
@@
Uncertainty estimation has shown promise in the identification of hallucinations in LLMs (Xiao & Wang, 2021; Huang et al., 2024b). High uncertainty often indicates the need for caution from users, as it suggests that the model is likely to be influenced by hallucinations (Zhang et al., 2023; Yoffe et al., 2024). In other words, high uncertainty signals unreliable responses. However, existing methods for uncertainty estimation in LLMs have limitations in estimating the inherent uncertainty of LLMs and lack efficiency. Including discussions of various heuristic methods, such as self-reflection (Ji et al., 2023), LLM uncertainty estimation can be categorized into sampling-based and token-based methods. Sampling-based methods estimate uncertainty by multiple sampling (Liu et al., 2024), perturbations (Zhang et al., 2024), or contrastive analysis (Huang et al., 2024a). Semantic Entropy (SE) (Kuhn et al., 2024) is a representative technique among these sampling-based methods. The core of these methods is estimating the consistency of LLM’s multiple guesses, and the response will be marked as unreliable if inconsistency is detected. However, sampling-based methods suffer from several limitations: (1) it cannot assess the reliability of a single response; (2) it requires multiple sampling iterations, making it inefficient and impractical for deployment in real-world applications; and (3) it fails to account for the model’s inherent uncertainty, such as consistently incorrect responses due to a lack of knowledge.
@@
Token-based methods can estimate the uncertainty of a single sentence without requiring multiple samplings as deterministic approaches (Gupta et al., 2024; Fadeeva et al., 2024). However, due to the lack of effective token-level uncertainty estimation, these methods often fail to achieve satisfactory results. For example, many tokens are correct but exhibit low probabilities (high entropy) in the model response. Many works have pointed out that the importance of different tokens in a sentence is not uniform (Lin et al., 2024a; Duan et al., 2024a), and the reliability of the model response depends mainly on a few critical tokens (Duan et al., 2024b; Bigelow et al., 2024), so they only focus on critical tokens when estimating reliability. However, as shown in Fig. 1(a), although both “Barack Hussein Obama” and “Abraham Lincoln” could be correct answers, their maximum probability of the critical token (first token of their names) is only $0.377$, which indicates a very high uncertainty. Therefore, to obtain an LLM uncertainty estimation that is both effective and efficient, a reliable token-level uncertainty estimation method is urgently required.
@@
In this paper, we find that the reason why the current token uncertainty estimation fails is that probability no longer captures the reliability of LLM responses. Specifically, probability can only reflect the relative strength relationship between different categories from the perspective of a discriminative model. However, generative LLMs are different from traditional discriminative models, and there may be more than one token that can be the suitable next-token. Therefore, measuring the probability of a single suitable candidate cannot capture the reliability of the model response. To capture the reliability of generating the next token, we need to assess whether the model has the knowledge to generate the next token, which can be referred to as the model’s inherent uncertainty. In this paper, we reveal that the strength of evidence accumulation is lost during normalization, leading to the failure of probability-based reliability estimation. Therefore, we present a new uncertainty estimation framework based on logits named Logits-induced Token Uncertainty (LogU), which achieves both effectiveness and efficiency. Specifically, LogU introduces a more clear token-level uncertainty estimation, decoupling uncertainty into tokens relative aleatoric uncertainty (AU) and model inherent epistemic uncertainty (EU). As illustrated in Fig. 2(a), distinct combinations of these two types of uncertainty are treated as separate cases. Compared to probability-based uncertainty modeling, LogU can not only express states of “I am sure” and “I do not know”, but it can also express “Lack knowledge, but I suggest” and “I know more than one answer”. Moreover, this framework enables the estimation of real-time reliability of any response without the need for sampling, ensuring high efficiency. To realize the LogU framework, we adapt evidence modeling (Sensoy et al., 2018) in this paper, treating logits as parameters of a Dirichlet distribution to characterize aleatoric uncertainty and epistemic uncertainty. Then, we validate the reliability of the estimated uncertainty by using it to guide downstream tasks. The main contributions of this paper can be summarized as: (1) We reveal why probability-based strategies fail and introduce a new framework to estimate the token uncertainty of LLMs, which shows that the token uncertainty is a promising way to estimate response reliability. (2) By employing evidential learning to estimate aleatoric uncertainty and epistemic uncertainty, we provide a viable implementation of the LogU framework. (3) We demonstrate two usage cases for the estimated uncertainty, thus validating its effectiveness and showcasing the significant potential of LogU.
(a) Logits-induced token uncertainty (LogU).
@@
Figure 2: Why LogU works? Left: Illustration of four different scenarios considered in LogU, where the gray bars represent the logits for predicting the next token, the triangular patterns represent the corresponding Dirichlet distribution, and the table below compares uncertainty estimation using probability with that using LogU. Right: A case study from a medical QA, where the markings under each word reflect reliability estimated according to LogU, as well as the values of AU (gray) and EU (blue). I: Both AU and EU are high, where LLaMA recommends a metal “Chromium” for diabetes patients. II: The total logits are low, but one token’s logit is larger than the others, indicating that the LLM lacks experience and knowledge but knows what should be the next token, where the LLM repeats the medicine “Glucomannan” that has been generated in the previous context. III: The LLM is very confident about the next token, where it generates the fixed phrase “has been”. IV: The LLM has enough knowledge and knows more than one suitable answer. For example, the LLM generates “[comma]”, which can be replaced by many other suitable words. The dilemma in Fig. 1 is addressed according to quadrant II and IV.
@@ 2 Related Work
Sampling-based uncertainty estimation. These methods evaluate the randomness in the LLM generation process. Specifically, they allow LLMs to guess multiple times and evaluate their consistency. (Cole et al., 2023) introduce repetition and diversity into the measurement of consistency. (Zhang et al., 2024) introduce a mechanism that perturbs semantically equivalent questions to evaluate the consistency of LLM responses across variants of the same question. (Huang et al., 2024a) inject correct and incorrect labels into the prompt during sampling, and the uncertainty level is defined based on whether the LLM’s responses remain consistent across three samplings for each instance. A more continuous measurement of uncertainty is based on similarity. A recent method proposed by (Lin et al., 2024b) calculates the similarity between multiple responses to indirectly quantify the dispersion of the model outputs. The most representative sampling-based method is Semantic Entropy (SE) (Kuhn et al., 2024), which improves token-level measures by clustering sentences into equivalence classes based on their semantic similarity and computing the entropy over these classes. However, sampling-based methods can not evaluate the model’s inherent uncertainty and are costly. Specifically, sampling-based methods only measure the consistency of multiple guesses but fail to identify whether the LLM lacks knowledge about the question.
Probability-based uncertainty estimation. Probability-based uncertainty estimation methods can also be described as deterministic methods. These methods calculate the confidence (uncertainty) of the model based on the probability distribution of the prediction (Arora et al., 2021). In addition to maximum probability, entropy is another common measure for estimating uncertainty. (Kadavath et al., 2022) use the probability of the complete sequence to compute predictive entropy for assessing the sharpness of the output distributions. However, not all tokens in a sentence are equally critical (Lin et al., 2024a). Recent work by (Duan et al., 2024a) highlights that not all tokens contribute equally to the underlying meaning, as linguistic redundancy often allows a few key tokens to capture the essence of longer sentences. However, even when focusing only on key tokens, these probability-based methods still cannot estimate the reliability of the answer generated by LLM. As shown in Fig. 1(a), the key tokens do not yet provide a reliable estimation of the risk.
@@ 3.2 Failures in Traditional Uncertainty Estimation
Traditional discriminative models typically use probability to estimate reliability. Probability, in fact, is a normalization of the strength of evidence for different categories. Due to the mutually exclusive properties of the different categories, the relative strength between different categories can accurately indicate the reliability of the prediction. However, in the case of LLMs, the situation is different. Although LLM’s next token prediction can still be viewed as a classification task with $|\bm{Y}|$ categories, these categories are no longer mutually exclusive. Even different tokens that are mutually exclusive in a conversation may no longer be mutually exclusive in a different context. This is why recent LLM research suggests shifting from token-to-token training to a concept-to-concept training paradigm. There may be more than one suitable next token, so the relative relationship between different tokens to no longer reflect the reliability of the response. Therefore, the information of the strength of the evidence before normalization becomes important. During the LLM training process, the suitable token accumulates evidence (logits increase, similar to the discriminative model (Wei et al., 2022)), specifically, the higher the strength of the evidence, the more similar scenarios the model has encountered during training.
@@
3.3 LogU: Four-quadrant Framework
The reason why probability-based methods fail to identify reliability is that probability is normalized. After logits are normalized, only the relative strength between different answers (should be either “Barack” or “George”) is retained, while the original strength information of the logits is discarded, which results in the loss of the ability to indicate reliability (distinguish between “I do not know” and “I know more than one answer”). To address this limitation, we propose a Logits-induced Token Uncertainty framework termed LogU. In addition to considering the relative relationships among tokens (AU), LogU also takes into account the strength of the model’s response (EU). With the information of EU, “I do not know” and “I know more than one answer” can be characterized separately. As shown in Fig. 2(a), the four quadrants of uncertainty are described as:
@@ 3.4 Considering Logits as Evidence
In this subsection, we present a viable implementation of the LogU framework, specifically by modeling AU and EU using a Dirichlet distribution. Inspired by Evidential Deep Learning (Sensoy et al., 2018), we treat logits as evidence for each token and model them into a Dirichlet distribution. A naive approach is using the logits of all non-negative values as evidence, while setting those with negative values to no evidence by applying the ReLU activation function. However, unlike conventional classification networks, the number of candidates generated by an LLM (that is, tokenizer size) is significantly large, with a considerable proportion of tokens having extremely low logits, and these tokens should be discarded as noise (Tang et al., 2024). Therefore, this paper focuses on the distribution of main candidates with high logits. Specifically, we sample the top $K$ tokens with the largest logits to model a Dirichlet distribution:
@@
and the underlying intuition is that larger $\alpha_{k}$ produces a sharper density, and thus it indicates increased confidence in a prediction. For more information on Dirichlet distribution, please refer to the introduction from (Ulmer et al., 2023).

### v3

@@ Abstract
Over the past few years, Large Language Models (LLMs) have developed rapidly and are widely applied in various domains. However, LLMs face the issue of hallucinations, generating responses that may be unreliable when the models lack relevant knowledge. To be aware of potential hallucinations, uncertainty estimation methods have been introduced, and most of them have confirmed that reliability lies in critical tokens. However, probability-based methods perform poorly in identifying token reliability, limiting their practical utility. In this paper, we reveal that the probability-based method fails to estimate token reliability due to the loss of evidence strength information which is accumulated in the training stage. Therefore, we present Logits-induced token uncertainty (LogTokU), a framework for estimating decoupled token uncertainty in LLMs, enabling real-time uncertainty estimation without requiring multiple sampling processes. We employ evidence modeling to implement LogTokU and use the estimated uncertainty to guide downstream tasks. The experimental results demonstrate that LogTokU has significant effectiveness and promise. Our code is available at link.
Index Terms:
Uncertainty estimation, Foundation models
@@ 1 Introduction
Over the past few years, Large Language Models (LLMs) have developed rapidly and LLM-driven systems are deployed across various domains. Despite their remarkable performance, LLMs remain prone to hallucinations [1], which causes them to generate unreliable responses when the models lack the corresponding knowledge. Hallucinations critically undermine the reliability of LLMs, particularly in professional applications such as medical and legal consultations [2, 3]. Hallucinations have been considered as a major barrier to the broader deployment of LLMs [4, 5, 6, 7].
@@
Uncertainty estimation has shown promise in the identification of hallucinations in LLMs [8, 9]. High uncertainty often indicates the need for caution from users, as it suggests that the model is likely to be influenced by hallucinations [10, 11]. In other words, high uncertainty signals unreliable responses. However, existing methods for uncertainty estimation in LLMs have limitations in estimating the inherent uncertainty of LLMs and lack efficiency. Including discussions of various heuristic methods, such as self-reflection [12], LLM uncertainty estimation can be categorized into sampling-based and token-based methods. Sampling-based methods estimate uncertainty by multiple sampling [5], perturbations [13], or contrastive analysis [14]. Semantic Entropy (SE) [15] is a representative technique among these sampling-based methods. The core of these methods is estimating the consistency of LLM’s multiple guesses, and the response will be marked as unreliable if inconsistency is detected. However, sampling-based methods suffer from several limitations: (1) it cannot assess the reliability of a single response; (2) it requires multiple sampling iterations, making it inefficient and impractical for deployment in real-world applications; and (3) it fails to account for the model’s inherent uncertainty, such as consistently incorrect responses due to a lack of knowledge.
@@
Token-based methods can estimate the uncertainty of a single sentence without requiring multiple samplings as deterministic approaches [16, 17]. However, due to the lack of effective token-level uncertainty estimation, these methods often fail to achieve satisfactory results. For example, many tokens are correct but exhibit low probabilities (high entropy) in the model response. Many works have pointed out that the importance of different tokens in a sentence is not uniform [18, 19], and the reliability of the model response depends mainly on a few critical tokens [20, 21], so they only focus on critical tokens when estimating reliability. However, as shown in Fig. 1(a), although both “Barack Hussein Obama” and “Abraham Lincoln” could be correct answers, their maximum probability of the critical token (first token of their names) is only $0.377$, which indicates a very high uncertainty. Therefore, to obtain an LLM uncertainty estimation that is both effective and efficient, a reliable token-level uncertainty estimation method is urgently required.
@@
In this paper, we find that the reason why the current token uncertainty estimation fails is that probability no longer captures the reliability of LLM responses. Specifically, probability can only reflect the relative strength relationship between different categories from the perspective of a discriminative model. However, generative LLMs are different from traditional discriminative models, and there may be more than one token that can be the suitable next-token. Therefore, measuring the probability of a single suitable candidate cannot capture the reliability of the model response. To capture the reliability of generating the next token, we need to assess whether the model has the knowledge to generate the next token, which can be referred to as the model’s inherent uncertainty. In this paper, we reveal that the strength of evidence accumulation is lost during normalization, leading to the failure of probability-based reliability estimation. Therefore, we present a new uncertainty estimation framework based on logits named Logits-induced Token Uncertainty (LogTokU), which achieves both effectiveness and efficiency. Specifically, LogTokU introduces a more clear token-level uncertainty estimation, decoupling uncertainty into tokens relative aleatoric uncertainty (AU) and model inherent epistemic uncertainty (EU). As illustrated in Fig. 2(a), distinct combinations of these two types of uncertainty are treated as separate cases. Compared to probability-based uncertainty modeling, LogTokU can not only express states of “I am sure” and “I do not know”, but it can also express “Lack knowledge, but I suggest” and “I know more than one answer”. Moreover, this framework enables the estimation of real-time reliability of any response without the need for sampling, ensuring high efficiency. To realize the LogTokU framework, we adapt evidence modeling [22] in this paper, treating logits as parameters of a Dirichlet distribution to characterize aleatoric uncertainty and epistemic uncertainty. Then, we validate the reliability of the estimated uncertainty by using it to guide downstream tasks. The main contributions of this paper can be summarized as: (1) We reveal why probability-based strategies fail and introduce a new framework to estimate the token uncertainty of LLMs, which shows that the token uncertainty is a promising way to estimate response reliability. (2) By employing evidential learning to estimate aleatoric uncertainty and epistemic uncertainty, we provide a viable implementation of the LogTokU framework. (3) We demonstrate two usage cases for the estimated uncertainty, thus validating its effectiveness and showcasing the significant potential of LogTokU.
(a) Logits-induced token uncertainty (LogTokU).
@@
Figure 2: Why LogTokU works? Left: Illustration of four different scenarios considered in LogTokU, where the gray bars represent the logits for predicting the next token, the triangular patterns represent the corresponding Dirichlet distribution, and the table below compares uncertainty estimation using probability with that using LogTokU. Right: A case study from a medical QA, where the markings under each word reflect reliability estimated according to LogTokU, as well as the values of AU (gray) and EU (blue). I: Both AU and EU are high, where LLaMA recommends a metal “Chromium” for diabetes patients. II: The total logits are low, but one token’s logit is larger than the others, indicating that the LLM lacks experience and knowledge but knows what should be the next token, where the LLM repeats the medicine “Glucomannan” that has been generated in the previous context. III: The LLM is very confident about the next token, where it generates the fixed phrase “has been”. IV: The LLM has enough knowledge and knows more than one suitable answer. For example, the LLM generates “[comma]”, which can be replaced by many other suitable words. The dilemma in Fig. 1 is addressed according to quadrant II and IV.
@@ 2 Related Work
Sampling-based uncertainty estimation. These methods evaluate the randomness in the LLM generation process. Specifically, they allow LLMs to guess multiple times and evaluate their consistency. [23] introduce repetition and diversity into the measurement of consistency. [13] introduce a mechanism that perturbs semantically equivalent questions to evaluate the consistency of LLM responses across variants of the same question. [14] inject correct and incorrect labels into the prompt during sampling, and the uncertainty level is defined based on whether the LLM’s responses remain consistent across three samplings for each instance. A more continuous measurement of uncertainty is based on similarity. A recent method proposed by [24] calculates the similarity between multiple responses to indirectly quantify the dispersion of the model outputs. The most representative sampling-based method is Semantic Entropy (SE) [15], which improves token-level measures by clustering sentences into equivalence classes based on their semantic similarity and computing the entropy over these classes. However, sampling-based methods can not evaluate the model’s inherent uncertainty and are costly. Specifically, sampling-based methods only measure the consistency of multiple guesses but fail to identify whether the LLM lacks knowledge about the question.
Probability-based uncertainty estimation. Probability-based uncertainty estimation methods can also be described as deterministic methods. These methods calculate the confidence (uncertainty) of the model based on the probability distribution of the prediction [25]. In addition to maximum probability, entropy is another common measure for estimating uncertainty. [26] use the probability of the complete sequence to compute predictive entropy for assessing the sharpness of the output distributions. However, not all tokens in a sentence are equally critical [18]. Recent work by [19] highlights that not all tokens contribute equally to the underlying meaning, as linguistic redundancy often allows a few key tokens to capture the essence of longer sentences. However, even when focusing only on key tokens, these probability-based methods still cannot estimate the reliability of the answer generated by LLM. As shown in Fig. 1(a), the key tokens do not yet provide a reliable estimation of the risk.
@@ 3.2 Failures in Traditional Uncertainty Estimation
Traditional discriminative models typically use probability to estimate reliability. Probability, in fact, is a normalization of the strength of evidence for different categories. Due to the mutually exclusive properties of the different categories, the relative strength between different categories can accurately indicate the reliability of the prediction. However, in the case of LLMs, the situation is different. Although LLM’s next token prediction can still be viewed as a classification task with $|\bm{Y}|$ categories, these categories are no longer mutually exclusive. Even different tokens that are mutually exclusive in a conversation may no longer be mutually exclusive in a different context. This is why recent LLM research suggests shifting from token-to-token training to a concept-to-concept training paradigm. There may be more than one suitable next token, so the relative relationship between different tokens to no longer reflect the reliability of the response. Therefore, the information of the strength of the evidence before normalization becomes important. During the LLM training process, the suitable token accumulates evidence (logits increase, similar to the discriminative model [27]), specifically, the higher the strength of the evidence, the more similar scenarios the model has encountered during training.
@@
3.3 LogTokU: Four-quadrant Framework
The reason why probability-based methods fail to identify reliability is that probability is normalized. After logits are normalized, only the relative strength between different answers (should be either “Barack” or “George”) is retained, while the original strength information of the logits is discarded, which results in the loss of the ability to indicate reliability (distinguish between “I do not know” and “I know more than one answer”). To address this limitation, we propose a Logits-induced Token Uncertainty framework termed LogTokU. In addition to considering the relative relationships among tokens (AU), LogTokU also takes into account the strength of the model’s response (EU). With the information of EU, “I do not know” and “I know more than one answer” can be characterized separately. As shown in Fig. 2(a), the four quadrants of uncertainty are described as:
@@ 3.4 Considering Logits as Evidence
In this subsection, we present a viable implementation of the LogTokU framework, specifically by modeling AU and EU using a Dirichlet distribution. Inspired by Evidential Deep Learning [22], we treat logits as evidence for each token and model them into a Dirichlet distribution. A naive approach is using the logits of all non-negative values as evidence, while setting those with negative values to no evidence by applying the ReLU activation function. However, unlike conventional classification networks, the number of candidates generated by an LLM (that is, tokenizer size) is significantly large, with a considerable proportion of tokens having extremely low logits, and these tokens should be discarded as noise [28]. Therefore, this paper focuses on the distribution of main candidates with high logits. Specifically, we sample the top $K$ tokens with the largest logits to model a Dirichlet distribution:
@@
and the underlying intuition is that larger $\alpha_{k}$ produces a sharper density, and thus it indicates increased confidence in a prediction. For more information on Dirichlet distribution, please refer to the introduction from [29].
3.5 Why Logits can Capture EU
Why evidential modeling can capture EU. Due to there is no ground-truth of the EU, uncertainty-aware model training (such as evidential deep learning [22]) does not obtain the EU through direct alignment. Instead, it is typically achieved by assuming that all scenarios should intuitively have a high EU for the model, and the EU is gradually eliminated during the model’s learning process. Therefore, there is a regularization constraint on the growth of the evidences (logits). This process consists of two parts: classification loss and regularization for constraining smaller logits. In other words, due to the trade-off between correct classification and logits regularization, logits are only allowed to be larger when the model is highly confident in its predictions, thereby characterizing the magnitude of the EU. Taking the training loss of typical evidential modeling ($\mathcal{L}_{\text{EVD}}$) as an example:
$\mathcal{L}_{\text{EVD}}=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Evidential deep learning classification loss}}\underbrace{+\mathcal{L}_{\text{reg}}}_{\text{logits constrainer}},$ (6)
where $z_{j}$ is the predictive logit of class $j$ and $\mathcal{L}_{\text{reg}}$ indicates the regularization for constraining the logits scale.
Cross-entropy can be interpreted as evidential modeling. Intuitively, the cross-entropy loss is not affected by the scale of logits; adding a constant to all elements of the logits vector does not change the probability values. For instance, both [20, 18] and [30, 28] are mapped to the probability distribution [0.88, 0.12] after applying the softmax function. However, intuition is not always trustworthy. During the actual gradient descent training of models, the scale of logits is constrained. In fact, as shown in Eq. 7, the cross-entropy loss can be decomposed into a format similar to the evidential deep learning loss function, which naturally penalizes shifting logits scales during gradient updates.
$\mathcal{L}_{\text{LLM}}$ $=-\sum_{j}^{|\bm{Y}|}y_{j}\log p_{j}$ (7)
$=-\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{e^{z_{j}}}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{z_{j^{\prime}}}}$
$=-\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{e^{(z_{j}+1})}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{(z_{j^{\prime}}+1})}$
$=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Evidential deeplearning classification loss}}$
$\quad-\sum_{j}^{|\bm{Y}|}y_{j}((z_{j}+1)-\log(z_{j}+1))$
$\quad\underbrace{-\log\frac{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{(z_{j^{\prime}}+1)}}}_{\text{logits constrainer}},$
In other words, absolute logits scale does not change probability distribution, but it affects optimization, which is the theoretical reason why logits can be considered as evidence similar to uncertainty-aware modeling.